ABC has $32 million of inventory and $31 million of accounts receivable. Its average daily sales are $851,000 and its gross profit margin is 22 percent. The company’s payables deferral period is 54.2 days. What is the length of the firm's cash conversion cycle?
Answer #1
Dear Student
Thank you for referring Chegg
Please give thumbs up if you are satisfied
Statement showing Computations
Particulars Amount
Cash conversion cycle = Days Inventory outstanding + days Sales o/S - Days Payable O/S
Daily Sales       851,000.00
Gross margin at 22%       187,220.00
COGS= Sales - Gross margin       663,780.00
Days Inventory outstanding = 32,000,000/663,780                  48.21
days Sales o/S = 31,000,000/851000                  36.43
Days Payable O/S =                    54.20
Cash conversion cycle = 48.21+36.43-54.20                  30.44
Dear Student
Thank you for referring Chegg
Please give thumbs up if you are satisfied
Statement showing Computations
Particulars Amount
Cash conversion cycle = Days Inventory outstanding + days Sales o/S - Days Payable O/S
Daily Sales       851,000.00
Gross margin at 22%       187,220.00
COGS= Sales - Gross margin       663,780.00
Days Inventory outstanding = 32,000,000/663,780                  48.21
days Sales o/S = 31,000,000/851000                  36.43
Days Payable O/S =                    54.20
Cash conversion cycle = 48.21+36.43-54.20                  30.44